OpenStudy (anonymous):

Sin2x/1+cos2x=4 find all exact values of x or to nearest degree

OpenStudy (mathmale):

which do you actually mean, 1) sin 2x / 1 + cos 2x =4, or 2) (sin 2x) / (1 + cos 2x) = 4 ? Those parentheses can make a huge difference in communication your intent accurately.

OpenStudy (aravindg):

Well it probably should be 2. No one wound write sin 2x/1 :)

OpenStudy (mathmale):

Better yet, consider using Equation Editor (below):\[\frac{ \sin 2x }{ 1+ \cos 2x }=4\]

OpenStudy (mathmale):

With this in mind, consider the double angle formulas for the sine and cosine. Hint: sin 2x = sin (x+x) = sin x cos x + cos x sin x = ?? Can you write a similar formula for cos 2x?

OpenStudy (mathmale):

\[\frac{ 2\sin x \cos x }{ 1 + formula~for~\cos 2x }=4\]